zarkam21:

help with 3 part question

zarkam21:

1 attachment
zarkam21:

1 attachment
Vocaloid:

for part a you would describe (with words) what the limitations for the square root expression are

Vocaloid:

something like, "the expression under the square root symbol must be 0 or greater"

Vocaloid:

then for b write an expression that describes what the possible x values are, given the previous condition

zarkam21:

okay so a would be " the expression under the square root symbol must be 0 or greater"

Vocaloid:

yes

zarkam21:

and for b the domain would be \[[2,\infty)\]

Vocaloid:

very nice

zarkam21:

\[, {x | x \ge 2}\]

zarkam21:

would i include this also, or just the first

Vocaloid:

either one is fine (since they're equivalent expressions)

zarkam21:

okay so for part II a

zarkam21:

i would use um an = something right

zarkam21:

i kind of remember

zarkam21:

a1-an maybe?

Vocaloid:



zarkam21:

an=4+(n-1)(3)

Vocaloid:

awesome so answer choice b)

zarkam21:

and a would just be 3 because that is the common difference

Vocaloid:

d is the common difference so d = 3

zarkam21:

Part III the two slopes are 3 and -2

zarkam21:

because the equations are in the form y=mx+b where m is the slope

Vocaloid:

yes

zarkam21:

okay and to determine the solutions , it has one right

Vocaloid:

good, that's it

zarkam21:

because they are different the slopes dont match

zarkam21:

Ugh thank you i appreciate all your help trust me